An electric field with a magnitude of 2.20 𝑘𝑁/𝐶 is applied along the x-axis. Calculate the



electric flux through a rectangular plane 0.77 m wide and 1.44 m long assuming that (a) the



plane is parallel to the yz plane; (b) the plane is parallel to the xy plane; (c) the plane



contains the y axis, and its normal makes an angle of 26.7° with the x axis.

1
Expert's answer
2022-02-24T09:04:54-0500

"\Phi=ES\cos\alpha,"

a)

"\alpha=0,"

"\Phi=ES=2.4~kVm,"

b)

"\alpha=90°,"

"\Phi=0,"

c)

"\alpha=26.7°,"

"\Phi=ES\cos26.7°=2.2~kVm."